Allison Corporation acquired 90 percent of Bretton on January 1, 2016. Of Bretton's total acquisition-date fair value, $60,000 was allocated to undervalued equipment (with a 10-year remaining life) and $80,000 was attributed to franchises (to be written off over a 20-year period).

Since the takeover, Bretton has transferred inventory to its parent as follows:

Year Cost Transfer Price Remaining at Year-End
2016 $ 45,000 $ 90,000 $ 30,000 (at transfer price)
2017 48,000 80,000 35,000 (at transfer price)
2018 69,000 92,000 50,000 (at transfer price)

On January 1, 2017, Allison sold Bretton a building for $50,000 that had originally cost $70,000 but had only a $30,000 book value at the date of transfer. The building is estimated to have a five-year remaining life (straight-line depreciation is used with no salvage value).

Selected figures from the December 31, 2018, trial balances of these two companies are as follows:

Allison Bretton
Sales $ 700,000 $ 400,000
Cost of goods sold 440,000 220,000
Operating expenses 120,000 80,000
Investment income Not given 0
Inventory 210,000 90,000
Equipment (net) 140,000 110,000
Buildings (net) 350,000 190,000

Determine consolidated totals for each of these account balances.
Answer #1

ANSWER

Consolidated Totals : Totals $1,008,000 $ 566,000 $ 206,000 Sales Cost of Goods sold Operating Expenses Investment Income Inv

Excess Amortization Expenses

Equipment $60,000/10 years = $6,000 per year

Franchises $80,000/20 years = 4,000 per year

Annual excess amortizations $10,000

Intra-entity Gross Profit—Inventory, 1/1/18:

Gross profit ($80,000 – $48,000) = $32,000

Gross profit rate ($32,000÷$80,000)= 40%

Remaining inventory = $35,000

Gross profit rate = 40%

Intra-entity Gross Profit in ending inventory, 1/1/18 (35000*40%) =14000

Intra-entity Gross Profit—Inventory, 12/31/18: Gross profit ($92,000 – $69,000) = $23,000

Gross profit rate ($23,000 ÷ $92,000) = 25%

Remaining inventory = $50,000

Gross profit rate = 25%

Intra-entity gross profit in ending inventory, 12/31/18 (50000*25%)=$12,500

Impact of Intra-Entity Building Transfer:

12/31/17—Transfer price figures Transfer price = $50,000

Gain on transfer ($50,000 – $30,000) = 20,000

Depreciation expense ($50,000 ÷ 5 years) = 10,000 Accumulated depreciation =10,000

12/31/18—Transfer price figures

Depreciation expense = 10,000

Accumulated depreciation = 20,000

12/31/17—Historical cost figures Historical cost = $70,000 Depreciation expense ($30,000 book value ÷ 5 years) = 6,000 Accumulated depreciation ($40,000 + $6,000) = 46,000

12/31/18—Historical cost figures

Depreciation expense =6,000

Accumulated depreciation = 52,000

CONSOLIDATED BALANCES

Sales = $1,008,000 (700000+400000-92000)

(add the two book values and subtract $92,000 in intra-entity transfers)

Cost of Goods Sold = $566,500 (440000+220000-92000-14000+12500)

(add the two book values and subtract $92,000 in intra-entity purchases. Subtract $14,000 because of the previous year deferred intra-entity gross profit and add $12,500 to defer the current year intra-entity gross profit in ending inventory.) Operating Expenses = $206,000 (120000+80000+10000-4000)

(add the two book values and include the $10,000 excess amortization expenses but remove the $4,000 in excess depreciation expense [$10,000 – $6,000] created by building transfer)

Investment Income = $0

(the intra-entity balance is removed because the individual revenue and expense accounts of the subsidiary are included for consolidation)

Inventory= $287,500 (210000+90000-12500)

(add the two book values and subtract the $12,500 ending intra-entity gross profit)

Equipment (net) = $292,000 (140000+110000+60000-18000)

(add the two book values and include the $60,000 allocation from the acquisition-date fair value less three years of excess amortizations)

Buildings (net) = $528,000 (350000+190000+8000-20000)

(add the two book values and subtract the $20,000 intra-entity gain on the transfer after two years of excess depreciation [$4,000 per year])

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************